How many grams of oxygen are formed when 6.21 mole
Answer #1

1) The balanced reaction is 4KO(s) + 2H2O(l) ---> 4KOH(s) + O2(g)

From the balanced reaction ,

4 moles of KOH formed along with 1 mole =32 g of O2

6.21 moles of KOH formed along with ( 6.21 mol * 32g) / 4 mole = 49.7 g of O2

2) The balanced reaction is 3NO2(g) + H2O(l) ---> 2HNO3(aq) + NO(g)

From the balanced reaction ,

2 moles = 2*63.02g = 126.04 g of HNO3 produced from 1 mole = 18.02 g of water

75.9 g of HNO3 produced from (75.9g*18.02g)/126.04g = 10.9 g of water

3)The balanced reaction is 2P+3Cl2--> 2PCl3

Molar mass(g/mole)          31     71        137.5

2 moles = 2*31=62 g of P reacts with 3 mole = 3*71=213 g of Cl2

12.39 g of P reacts with (12.39*213)/62 = 42.54 g of Cl2

So both reactants all completely reacted

So 2 mole = 2*31=62 g of P produces 2 mole = 2*137.5 g of PCl3

12.39 g of P produces (12.39*2*137.5)/62 = 54.93 g of PCl3

(4) The balanced reaction is BCl3 + 3H2O ---> H3BO3 + 3HCl

Molar mass ( g/mole)          117          18                         36.5

1 mole = 117 g of BCl3 reacts with 3 moles = 3*18 = 54 g of H2O

60.0 g of BCl3 reacts with (60.0g*54g)/117g = 27.7 g of water

So (37.5 - 27.7)g of water left unreacted and is the excess reactant

Since all the mass of BCl3 reacted completely BCl3 is the limiting reactant

From the balanced reaction ,

1 mole = 117 g of BCl3 produces 3 moles = 3*36.5g of HCl

60.0 g of BCl3 produces ( 60.0g*3*36.5g) / 117g = 56.0 g of HCl